# Label Anything: Multi-Class Few-Shot Semantic Segmentation with Visual Prompts

### v1

@@ 4.4 Ablation Study
Our comprehensive ablation studies on the validation fold of COCO-$20^{i}$ dissected the contribution of each component within our model to its overall performance. We systematically omitted elements deemed non-critical—specifically, the Spatial Convolutions in the Mask Decoder, the Class Token Pool, and the Class Example Mixer from the Prompt Encoder—and observed the resultant impact on model efficacy. As delineated in Table 4, excluding the Spatial Convolutions from the Mask Decoder markedly diminished model performance. This component is pivotal for blending spatial details before classification, significantly enhancing segmentation quality. Meanwhile, the Class Token Pool and Class Example Mixer, though their removal slightly affected performance, were proven indispensable for optimizing results. These findings highlight the integral role of each component in achieving superior segmentation outcomes, underscoring the holistic synergy that drive our model’s performance.
Our analysis also assessed the impact of different prompt types on our model’s effectiveness, exploring how each type influences performance. Table 5 reveals that mask prompts led to the highest performance, offering the most detailed guidance for segmentation tasks. Nonetheless, the model demonstrated robust adaptability, yielding solid results even with alternative prompt types, albeit with a marginal reduction compared to masks alone. This underscores the model’s versatile capability to effectively manage and interpret various prompt types.
@@
w/o Class Example Mixer 38.7 30.9
@@
Table 4: Ablation analysis on the COCO-$20^{i}$’s fold-0 evaluating the impact of different components of the Label Anything architecture.
@@
Table 5: Ablation analysis evaluating the impact of different prompt types on model performance.
5 Conclusion
In this study, we presented Label Anything, an innovative framework for few-shot semantic segmentation that markedly advances the adaptability and efficiency of segmentation models. Our rigorous testing on the COCO-$20^{i}$ benchmark has showcased the model’s capacity to reach competitive and, in some instances, state-of-the-art performance, alongside its ability to generalize across diverse segmentation scenarios. Including a wide range of visual prompts—points, bounding boxes, and masks—empowers Label Anything to interpret and leverage multiple forms of input guidance effectively. The results highlight the efficacy of prototype-based approaches in addressing the challenges of multi-class FSS, presenting a compelling alternative to affinity-based methods which, despite their accuracy, struggle with scalability as the dataset expands.
We envision a broad spectrum of future research directions to advance our method. An intriguing avenue involves the integration of textual prompts, leveraging the interplay between textual and visual information to enrich the segmentation process. Concurrently, the proposal of an optimized example candidate selection mechanism, potentially through clustering approaches, promises to streamline the identification of example candidates within large-scale datasets, enhancing the efficiency of support set construction.
5.0.1 Acknowledgement
We acknowledge the CINECA award under the ISCRA initiative, which gave us access to computing resources and support.

### v2

@@ 5.4 Ablation Study
Our comprehensive ablation studies on the validation fold of COCO-$20^{i}$ examined the contribution of each component within our model to its overall performance. We systematically omitted non-essential elements—specifically, the spatial convolutions in the mask decoder, the class token pool, and the class-example mixer from the prototype learner—and observed the resultant impact on model efficacy. As delineated in Table 5, excluding the spatial convolutions from the mask decoder markedly diminished model performance. This component is pivotal for blending spatial details before classification, significantly enhancing segmentation quality. Meanwhile, the class token pool and class-example mixer, though their removal slightly affected performance, were proven indispensable for optimizing results. These findings highlight the integral role of each component in achieving superior segmentation outcomes, underscoring the holistic synergy that drives the performance of our model. Our analysis also assessed the impact of different prompt types on the effectiveness of our model, exploring how each type influences its performance. Table 6 reveals that mask prompts led to the highest performance, offering the most detailed guidance for segmentation tasks. Nonetheless, the model demonstrated robust adaptability, yielding solid results even with alternative prompt types, albeit with a marginal reduction compared to masks alone. This underscores the model’s versatile capability to manage and interpret various prompt types effectively.
6 Conclusion
In this study, we presented Label Anything, an innovative framework for few-shot semantic segmentation that advances the adaptability and efficiency of segmentation models. Our rigorous testing on the COCO-$20^{i}$ benchmark has demonstrated the model’s capacity to achieve competitive and, in some instances, state-of-the-art performance, while also showcasing its ability to generalize across diverse segmentation scenarios. Including a wide range of visual prompts—points, bounding boxes, and masks—empowers Label Anything to effectively interpret and leverage multiple forms of input guidance. The results highlight the efficacy of prototype-based approaches in addressing the challenges of multi-class FSS, presenting a compelling alternative to affinity-based methods, which, despite their accuracy, struggle with scalability as the dataset expands.
We envision several future research directions to develop our method further. An intriguing avenue involves a prototype-affinity multi-class hybrid method to achieve both efficiency and accuracy. This approach, coupled with the development of algorithms for auxiliary image search through similarity-based methods, aims to augment the support set with highly relevant images, thereby enhancing the model’s adaptability to diverse scenarios. Concurrently, the proposal of an optimized example candidate selection mechanism promises to streamline candidate identification within large-scale datasets.
Table 5: Ablation analysis on the COCO-$20^{i}$’s fold-0 evaluating the impact of different components of the LA architecture.
@@ 6 Conclusion
w/o Class-example mixer 38.7 30.9
@@
Table 6: Ablation analysis on the impact of different prompt types on model performance.
@@
{ack}
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).